Mary is 3 times as old as Ben. After four years, their total age is 41 years old. Find the age, in years and months, of Mary

Answer by addingup(3677)

M = 3B
M+4+B+4 = 41 substitute for M:
3B+4+B+4 = 41 consolidate similar terms
3B+B+4+4 = 41
4B+8 = 41
4B = 33
B = 33/4 = 8.25
M = 3B
M = 3(8.25) = 24.75